OpenStudy (anonymous):

Bank A offers a checking account costing $ 3 per month plus $0.20 per check after the first 10 checks. Bank B offers Unlimited Checking for a $6 monthly Charge. A) Which Bank is Cheaper if You write 20 Checks per month? B) How many checks do you have to write so the cost is the same in both banks?

OpenStudy (ganpat):

Bank A ; 3 + 0.20(x - 10).. where x = no of checks .. Bank B ; 6$.. part A , 20 checks.. Bank A; 3 + 0.2 (20 -10) = 3 + 2 = 5 $ Bank B; 6$ So which one is cheaper ?? Part B , Number of checks ?? equate 2 equations.. 3 + 0.2(x - 10) = 6 0.2x - 2 = 3 0.2x = 5 x = 25 checks..

OpenStudy (anonymous):

I get it now, so basically your just solving for part A. I appreciate the assistance.

OpenStudy (ganpat):

i solved for both.. Part A & Part B...

OpenStudy (anonymous):

oh yeah